# Revisiting End-to-End Learning with Slide-level Supervision in Computational Pathology

### v1

@@ Appendix A Theoretical Analysis of Optimization Risk in End-to-End Training
A.1 Multi-head Local Attention Mechanism
@@
$\Rightarrow\mathcal{R}_{\text{MHLA}}=\sum_{j}^{m}\mathcal{R}^{(j)}<\underbrace{m\cdot\max_{1\leq j\leq m}\mathcal{R}^{(j)}}_{\color[rgb]{0.22265625,0.7109375,0.2890625}\text{upper bound of multi-head risk}}=m\cdot\underbrace{\frac{1}{m}\mathcal{R}_{\text{ABMIL}}}_{\color[rgb]{0.22265625,0.7109375,0.2890625}{\text{each head have 1/m impact}}}$ (9)
@@
A.2 Attention Plus Propagation
@@ B.2 Preprocess
End-to-End Training. To efficiently process gigapixel slides in our E2E pipeline, we first crop each WSI into a series of non-overlapping patches of size (256 × 256) and discard background regions, including holes, as in CLAM [38]. Except for PANDA, we perform patching at 10x magnification, with an average of approximately 3,000 patches per slide. For the PANDA dataset, we process patches at 40x magnification, resulting in an average of 505 patches. Additionally, we extract patches at 5x and 20x magnifications to support our multi-scale random sampling strategy. After patch extraction, we store all data using the LMDB (Lightning Memory-Mapped Database) format. Ultimately, each dataset contains an average of 4$\sim$6 million 256×256 patches.
Two-Stage Framework. Following prior works [38, 46, 62, 54], we crop each WSI into a series of non-overlapping patches of size (256 $\times$ 256) at 20$\times$ magnification and discard the background regions, including holes, as in CLAM [38]. The average number of patches per dataset is around 10,000. To efficiently handle the large number of patches, we follow the traditional two-stage paradigm, using a pre-trained offline model to extract patch features. This includes a ResNet-50 [21] pre-trained on ImageNet-1k [15]. Specifically, the last convolutional module of the ResNet-50 is removed, and a global average pooling is applied to the final feature maps to generate the initial feature vector. Additionally, we also use state-of-the-art foundation models pre-trained on WSIs, such as CHIEF [58], GigaPath [60] and UNI [10].
@@ B.3 Implementation Details
End-to-End Training. To maintain consistency with traditional ResNet-based two-stage methods, we removed the last stage module of ResNet. For MIL, we added a LayerNorm [1] layer at its input to better optimize the Encoder. We also disabled the bias of all fully connected layers in MIL, which we found beneficial for E2E optimization. For training, we used different hyperparameters for different tasks. For cancer grading (PANDA), we employed an Adam [28] optimizer with a learning rate of $2\times 10^{-4}$ and a weight decay of $1\times 10^{-5}$, training for 200 epochs. For sub-typing (NSCLC, BRCA), we used an AdamW [37] optimizer with a learning rate of $8\times 10^{-5}$ and no weight decay, training for 75 epochs. For survival analysis (LUAD, BLCA, BRCA), we utilized an AdamW optimizer with a learning rate of $8\times 10^{-5}$ and a weight decay of $5\times 10^{-2}$, training for 30 epochs. The learning rate was adjusted using the Cosine annealing strategy. During training, we applied simple geometric data augmentations such as flipping and RandomResizedCrop. Since slides are typically H&E-stained, we found that color-related data augmentations could significantly impact performance. All experiments are conducted on 3090 GPUs. We adjusted the batch size based on the 24GB memory limit and the number of samples in different datasets. Unless otherwise specified, all ablation experiments in the manuscripts and supplementary materials use ResNet-18 as the encoder. All efficiency experiments are conducted on the BRCA-subtyping benchmark. We calculate the FLOPs for all models using an input size of 1$\times$512$\times$3$\times$224$\times$224, which simulates the sampling of 512 patches during E2E training. To evaluate model inference speed, we use an input size of 1$\times$10000$\times$3$\times$224$\times$224, representing the average data volume processed in clinical scenarios.
@@ C.1 More about E2E Methods
Slide-level supervised E2E methods process entire slides in a unified manner, preserving the critical interdependencies necessary for robust clinical interpretation. These methods exploit the complete spatial context of gigapixel images, thereby providing a more comprehensive and clinically pertinent analysis. Building on these merits, several researchers have introduced innovative slide-level supervised E2E methods to further enhance WSI analysis. Sharma et al. [48] proposed an E2E framework (C2C) that clusters patch representations and employs adaptive attention with KL-divergence regularization to robustly classify whole slide images. Li et al. [30] proposed a task-specific fine-tuning framework (FT) that employs a variational information bottleneck to distill patches into a sparse subset via Monte Carlo-sampled Bernoulli masks, thereby enabling E2E backbone fine-tuning.
@@ Appendix E Additional Related Works
As elaborated in the Related Work section, current E2E methods can be divided into instance-level supervised approaches and slide-level supervised approaches. Instance-level supervised methods [12, 39, 44, 45] adopt a pseudo E2E paradigm, in which the encoder is trained using instance-level pseudo-labels rather than genuine slide-level supervision. This strategy simplifies the problem by processing patches independently; however, it neglects the essential inter-patch contextual relationships required for robust clinical interpretation [6, 40] and creates a disconnect between training and downstream clinical applications. Moreover, the performance of these methods fundamentally depends on the quality of the pseudo-labels, indicating that they constitute a compromise rather than a fully E2E solution. For example, Qu et al. [45] introduce a novel instance-level MIL framework that leverages weakly supervised contrastive learning and prototype-based pseudo label generation to markedly improve both instance and bag-level classification in WSI analysis. Luo et al. [39] propose a negative instance guided self-distillation framework that leverages true negative samples and a prediction bank to constrain pseudo-label distributions, enabling an E2E instance-level classifier.

### v2

@@ A.1 Definition of Optimization Risk
A.2 Multi-head Local Attention Mechanism
@@
[…]
@@
A.3 Attention Plus Propagation
@@
A.4 Empirical Validation of Theoretical Analysis
Our theoretical analysis posits that the maximum attention score of noisy instances serves as a proxy for optimization risk. The theory further suggests that ABMILX can mitigate this risk while maintaining reasonable sparsity. To empirically validate this connection, we introduce the MAX-N metric, defined as the product of the maximum attention score of noisy instances for each slide and the total number of instances.
We measured MAX-N on the CAMELYON dataset during the early E2E training stage, with the results reported in Table 5. The experimental results align closely with our theoretical analysis. ABMILX drastically reduces the MAX-N score (our risk proxy) from 21.2162 (for ABMIL) to 2.6557, demonstrating its effectiveness in mitigating optimization risk. Concurrently, it maintains a functional sparsity level (36) compared to the baseline’s (80), achieving the "reasonable sparsity" predicted by our theory. This effective risk mitigation and balanced sparsity directly translate to a substantial performance improvement (95.88% vs. 91.78%). These findings validate our theoretical framework, demonstrating a clear alignment between our analysis and the experimental results.
Metric/Method ABMIL ABMILX
MAX-N 21.2162 2.6557
Sparsity 80 36
Performance 91.78 95.88
Table 5: Empirical validation of ABMILX’s risk mitigation on the CAMELYON dataset. ABMILX significantly reduces the optimization risk proxy (MAX-N) and balances sparsity, aligning with our theoretical analysis and leading to superior performance over the ABMIL baseline.
@@ B.2 Preprocess
[…]
@@ B.3 Implementation Details
[…]
@@ C.1 More about E2E Methods
Slide-level supervised E2E methods process entire slides in a unified manner, preserving the critical interdependencies necessary for robust clinical interpretation. These methods exploit the complete spatial context of gigapixel images, thereby providing a more comprehensive and clinically pertinent analysis. Building on these merits, several researchers have introduced innovative slide-level supervised E2E methods to further enhance WSI analysis. Sharma et al. [50] proposed an E2E framework (C2C) that clusters patch representations and employs adaptive attention with KL‐divergence regularization to robustly classify whole slide images. Li et al. [30] proposed a task-specific fine-tuning framework (FT) that employs a variational information bottleneck to distill patches into a sparse subset via Monte Carlo-sampled Bernoulli masks, thereby enabling E2E backbone fine-tuning.
@@ Appendix E Additional Related Works
While E2E high-resolution image analysis is relatively mature in general computer vision [45, 44, 5], its application to CPath presents significant challenges due to their unique gigapixel scale. As elaborated in the Related Work section, current E2E methods for WSI analysis are broadly categorized into instance-level supervised and slide-level supervised approaches. As elaborated in the Related Work section, current E2E methods can be divided into instance-level supervised approaches and slide-level supervised approaches. Instance-level supervised methods [13, 39, 46, 47] adopt a pseudo E2E paradigm, in which the encoder is trained using instance-level pseudo‐labels rather than genuine slide-level supervision. This strategy simplifies the problem by processing patches independently; however, it neglects the essential inter-patch contextual relationships required for robust clinical interpretation [7, 40] and creates a disconnect between training and downstream clinical applications. Moreover, the performance of these methods fundamentally depends on the quality of the pseudo‐labels, indicating that they constitute a compromise rather than a fully E2E solution. For example, Qu et al. [47] introduce a novel instance-level MIL framework that leverages weakly supervised contrastive learning and prototype-based pseudo label generation to markedly improve both instance and bag-level classification in WSI analysis. Luo et al. [39] propose a negative instance guided self-distillation framework that leverages true negative samples and a prediction bank to constrain pseudo-label distributions, enabling an E2E instance-level classifier.